Answer each of the following questions. You have to show all your work to get full e Problema 1. (5 marks; 3, 2) Assume X ~ Gamma(ai, β) and Y ~ Gamma(a2, β) are independent random variables. a) Compute the joint density of U = X X+Y and VX/X+Y), be sure to include the support/domain.
Answer #1

Problem 1 The joint distribution of X and Y is fxy(x,y) = since X and Y are independent and r,y >0 「(ai)「(a2) 0 otherwise Tra

\begin{vmatrix} J\left(\frac{x,y}{u,v} \right ) \end{vmatrix}=\begin{vmatrix} \frac{\partial x}{\partial u} & \frac{\partial y}{\partial u}\\ \frac{\partial x}{\partial v} & \frac{\partial y}{\partial v} \end{vmatrix}=\begin{vmatrix} v &1-v \\ u& -u \end{vmatrix}=-u,~|Jacobian|=u\\ The~joint~distribution~of~U,~V~is\\ f_{U,V}(u,v)=\frac{\beta^{\alpha_1+\alpha_2}u^{\alpha_1+\alpha_2-1}v^{\alpha_1-1}(1-v)^{\alpha_2-1}\exp(-\beta u)}{\Gamma(\alpha_1)\Gamma(\alpha_2)}\\ =\left(\frac{\beta^{\alpha_1+\alpha_2}}{\Gamma(\alpha_1+\alpha_2)} \exp(-\beta u)u^{\alpha_1+\alpha_2-1}\right )\left(\frac{1}{B(\alpha_1,\alpha_2)}v^{\alpha_1-1}(1-v)^{\alpha_2-1} \right ),\\ u>0,~0<v<1\\ =0~otherwise\\ Hence~U~and`V~are~independently~distributed~where\\ U\sim Gamma(\alpha_1+\alpha_2,\beta)~and~V\sim Beta(\alpha_1,\alpha_2).